# Equal contribution    Project lead    Corresponding author

### v1

@@ 1 Introduction
    We introduce Time-MoE, a universal, decoder-only time series forecasting foundation model architecture with mixture-of-experts, scaling up to 2.4 billion parameters for the first time. Time-MoE achieves substantial improvements in forecasting precision and consistently outperforms dense models with similar computational budgets, all while maintaining high efficiency.
@@
    Trained on Time-300B, Time-MoE models outperform other time series foundation models with a similar number of activated parameters across six real-world benchmarks, achieving reductions in forecasting errors by an average of 23% and 25% in zero-shot and in-distribution scenarios, respectively.
@@ 3 Methodology
Problem Statement. We address the problem of predicting future values in a time series: given a sequence of historical observations $\mathbf{X}_{1:T}=\left(x_{1},x_{2},\ldots,x_{T}\right)\in\mathbb{R}^{T}$ spanning $T$ time steps, our objective is to forecast the next $H$ time steps, i.e., $\hat{\mathbf{X}}_{T+1:T+H}=f_{\theta}\left(\mathbf{X}_{1:T}\right)\in\mathbb{R}^{H}$. Here, $f_{\theta}$ represents a time series model, where $T$ is the context length and $H$ is the forecasting horizon. Notably, both $T$ and $H$ are arbitrary when inferencing Time-MoE, distinguishing it from task-specific models with fixed configurations. Additionally, channel independence (Nie et al., 2023) is adopted to transform a multivariate input into univariate series, allowing Time-MoE to handle any-variate forecasting problems in real-world applications.
Figure 2: The architecture of Time-MoE, which is a decoder-only model. Given an input time series of arbitrary length, 1 we first tokenize it into a sequence of data points, 2 which are then encoded. These tokens are processed through $N$-stacked backbone layers, primarily consisting of causal multi-head self-attention and 3 sparse temporal mixture-of-expert layers. During training, 4 we optimize forecasting heads at multiple resolutions. For model inference, Time-MoE provides forecasts of arbitrary length by 5 dynamically scheduling these heads.
@@ 3.1 Time-MoE Overview
$g_{N+1,t}$ $=\operatorname{Sigmoid}\left(\mathbf{W}^{l}_{N}\mathbf{\bar{u}}^{l}_{t}\right),$ (7)
@@
Multi-resolution Forecasting. We introduce a novel multi-resolution forecasting head, which allows for forecasting at multiple scales simultaneously, in contrast to existing foundation models that are limited to a single fixed scale. This capability enhances Time-MoE ’s flexibility by enabling it to generate forecasts across various horizons. The model uses multiple output projections from single-layer FFNs, each designed for different prediction horizons. By incorporating a simple greedy scheduling algorithm (see Appendix B), Time-MoE efficiently handles predictions across arbitrary horizons. This design also boosts prediction robustness through multi-resolution ensemble learning. Additionally, Time-MoE aggregates forecasting errors from different horizons to compute a composite loss (Section 3.2.2), thereby improving the model generalization.
@@ 3.2.1 Time-300B Dataset
Training time series foundation models requires extensive, high-quality data. However, large-scale datasets that are adequately processed remain scarce. Recent advancements have facilitated the collection of numerous time series datasets from various sources (Godahewa et al., 2021; Ansari et al., 2024; Woo et al., 2024; Liu et al., 2024b). Nonetheless, data quality remains a challenge, with prevalent issues such as missing values and invalid observations (Wang et al., 2024a) that can impair model performance and destabilize training. To mitigate these issues, we developed a streamlined data-cleaning pipeline (Appendix C) to filter and refine raw data, and constructed the largest open-access, high-quality time series data collection named Time-300B for model pre-training. Time-300B is composed of a diverse range of publicly available datasets, spanning multiple domains such as energy, retail, healthcare, weather, finance, transportation, and web, as well as a portion of synthetic data to enhance the data quantity and diversity. Time-300B covers a wide spectrum of sampling frequencies, spanning from seconds to yearly intervals, and has over 300 billion time points after being processed by our data-cleaning pipeline, as summarized in Table 1.
@@
3.2.2 Loss Function
Pre-training time series foundation models in large scale presents significant challenges in training stability due to the massive datasets and the vast number of parameters involved. To address this, we use the Huber loss (Huber, 1992; Wen et al., 2019), which provides greater robustness to outliers and improves training stability.
where $\delta$ is a hyperparameter that balances the L1 and L2 loss components. When training the model with a MoE architecture, focusing solely on optimizing prediction error often leads to load imbalance issues among the experts. A common problem is routing collapse (Shazeer et al., 2017), where the model predominantly selects only a few experts, limiting training opportunities for others. To mitigate this, following the approaches of (Dai et al., 2024; Fedus et al., 2022), we achieve expert-level balancing with an auxiliary loss to reduce routing collapse:
$\mathcal{L}=\frac{1}{P}\sum_{i=1}^{P}\mathcal{L}_{\text{ar}}\left(\mathbf{X}_{t+1:t+p_{i}},\hat{\mathbf{X}}_{t+1:t+p_{i}}\right)+\alpha\mathcal{L}_{\text{aux}},$ (11)
3.2.3 Model Configurations and Training Details
[…]
@@ 4 Main Results
Time-MoE consistently outperforms state-of-the-art forecasting models by large margins across six well-established benchmarks and settings (Appendix B). We compared Time-MoE against a wide range of time series models, including several recently released foundation models for forecasting. To ensure a fair comparison, we adhered to the experimental configurations from (Woo et al., 2024) for out-of-distribution forecasting and (Wu et al., 2023) for in-distribution forecasting, applying a unified evaluation pipeline that we developed.
Baselines. We evaluate Time-MoE against 14 different baselines, representing state-of-the-art models in long-term forecasting. These baselines are categorized into two groups. The first group, for zero-shot forecasting evaluation, includes pre-trained foundation models such as Moirai (2024), TimesFM (2024), Timer (2024b), Moment (2024), and Chronos (2024). The second group, for in-distribution (full-shot) forecasting evaluation, consists of deep time series models such as iTransformer (2024a), TimeMixer (2024b), TimesNet (2023), PatchTST (2023), Crossformer (2023), TiDE (2023), DLinear (2023), and FEDformer (2022b).
@@ 4.1 Zero-shot Forecasting
Table 3: Full results of zero-shot forecasting experiments. A lower MSE or MAE indicates a better prediction. TimesFM, due to its use of Weather and Electricity datasets in pretraining, is not evaluated on these two datasets and is denoted by a dash (--). Red: the best, Blue: the 2nd best.
@@
Results. Detailed results of zero-shot forecasting are in Table 3. Time-MoE achieves consistent state-of-the-art performances, improving a large margin as MSE reduction in average exceeding 23% over the other most competitive baselines. Importantly, as the model size scales (e.g., base $\rightarrow$ ultra), it continuously exhibits enhanced performance across all datasets, affirming the efficacy of scaling laws within our time series foundation models. Furthermore, in comparisons with robust baselines possessing a similar number of activated parameters, Time-MoE demonstrates significantly superior performance. Compared to Moirailarge and Chronoslarge, Time-MoE achieved average MSE reductions of 12% and 24% respectively, and compared to Moment, it achieved a relative improvement of 41%.
@@ 4.2 In-distribution Forecasting
Table 4: Full results of in-domain forecasting experiments. A lower MSE or MAE indicates a better prediction. Full-shot results are obtained from (Liu et al., 2024a). Red: the best, Blue: the 2nd best.
@@
0.286 0.336 0.285 0.333 0.281 0.332 0.328 0.357 0.315 0.350 0.335 0.377 0.333 0.362 0.540 0.493 0.408 0.424 0.374 0.398 0.358 0.399
$1^{\text{st}}$ Count 2 13 20 0 5 0 0 0 0 0 0
Setup. We fine-tune the pre-trained Time-MoE models on the train split of the above-mentioned six benchmarks and set the number of training epochs to only one.
Results. The full results are in Table 4. Time-MoE exhibits remarkable capabilities, comprehensively surpassing advanced deep time series models from recent years, achieving an average MSE reduction of 25%. Fine-tuning on downstream data with only one epoch significantly improves predictive performance, showcasing the remarkable potential of large time series models built on the MoE architecture. Similar to zero-shot forecasting, as the model size increases, the scaling law continues to be effective, leading to continuous improvements in the performance of the Time-MoE.
@@ 4.3 Ablation Study
Time-MoEbase 0.245
         w/o Huber loss 0.250
         w/o multi-resolution layer 0.252
         w/o mixture-of-experts 0.255
         w/o auxiliary loss 0.258
@@
Time-MoEbase 0.245 0.095 s/iter
Time-MoEbase w/ {1,8,32} 0.256 0.130 s/iter
Time-MoEbase w/ {1,8} 0.302 0.411 s/iter
Time-MoEbase w/ {1} 1.346 2.834 s/iter
@@
Replacing the MoE layers with standard FFNs (w/o mixture-of-experts) led to an average performance drop from $0.245$ to $0.255$, highlighting the performance boost provided by the sparse architecture. A detailed comparison of dense and sparse models is presented in Section  4.4. We retained only the horizon-32 output layer by eliminating the other multi-resolution output layers from the Time-MoEbase, excluding the multi-task optimization (w/o multi-resolution layer). Consequently, we observed that the performance of this modified model was slightly inferior compared to that of the Time-MoEbase. Additionally, as shown in the right side of Table 5, our default selection of four multi-resolution output projections with receptive horizons of $\{1,8,32,64\}$ results in optimal predictive performance and inference speed. As we reduce the number of multi-resolution output projections, performance consistently declines, and inference speed significantly increases. This demonstrates the rationality of our multi-resolution output projection design.
@@ 4.4 Scalability Analysis
. Average MSE Training Speed Inference Speed Training Memory Inference Memory Time-MoEbase 0.245 0.84 s/iter 0.095 s/iter 1.77 GB 226.70 MB Time-MoEbase w/o FA 0.245 1.09 s/iter 0.118 s/iter 1.77 GB 226.70 MB Time-MoEbase w/ FP32 0.244 1.24 s/iter 0.133 s/iter 2.21 GB 453.41 MB
@@ 4.5 Sparsification Analysis
w/ {Top1} 0.246 0.082 s/iter
w/ {Top2} 0.245 0.095 s/iter
w/ {Top4} 0.245 0.109 s/iter
w/ {Top6} 0.246 0.120 s/iter
w/ {Top8} 0.248 0.129 s/iter
We also performed a sensitivity analysis on the number of experts, represented as topk, within the MoE architecture, as shown in Table 7. As $k$ increases, performance shows only marginal changes, with minimal improvements in average MSE. However, inference time increases noticeably as more experts are utilized. This indicates that increasing sparsity within the MoE architecture does not compromise performance but significantly enhances computational efficiency. This balance is critical for scaling large models, where optimizing performance and computational cost is essential. Sparse MoE architectures inherently offer advantages in these areas.
@@ 5 Conclusion
In this paper, we introduced Time-MoE, a scalable and unified architecture for time series foundation models that leverages a sparse design with mixture-of-experts to enhance computational efficiency without compromising model capacity. Pre-trained on our newly introduced large-scale time series dataset, Time-300B, Time-MoE was scaled to 2.4 billion parameters, with 1.1 billion activated, demonstrating significant improvements in forecasting accuracy. Our results validate the scaling properties in time series forecasting, showing that Time-MoE consistently outperforms dense models with equivalent computational budgets across multiple benchmarks. With its ability to perform universal forecasting and superior performance in both zero-shot and fine-tuned scenarios, Time-MoE establishes itself as a state-of-the-art solution for real-world forecasting challenges. This work sets the stage for future advancements in scaling and developing more efficient time series foundation models.
@@ Appendix B Implementation Details
We evaluate the performance of various models for long-term forecasting across eight well-established datasets, including the Weather, Electricity, and ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2). A detailed description of each dataset is provided in Table 9.
@@
To construct the multi-resolution forecasting head, we define $P$ output projections, each corresponding to a distinct forecasting horizon, denoted as $\left(p_{1},p_{2},\ldots,p_{P}\right)$. The output projection for horizon $p_{i}$ is used to forecast the subsequent $p_{i}$ time steps, as follows:
$\hat{\mathbf{X}}_{t+1:t+p_{i}}=\mathbf{W}_{p_{i}}\mathbf{h}^{L}_{t},$ (12)
where $\mathbf{W}_{p_{i}}\in\mathbb{R}^{p_{i}\times D}$ is the learnable parameter matrix for that horizon, and $\mathbf{h}^{L}_{t}$ represents the output hidden state from the last MoE Transformer block. All output projections are optimized simultaneously during model training.
During inference, we apply a greedy scheduling algorithm for arbitrary target output lengths $H$, as outlined in Algorithm 1. For each forecast operation in the auto-regressive process, we select a projection $p_{i}$ with the closest forecasting horizon that does not exceed the remaining forecast duration. This approach allows Time-MoE to extend predictions beyond the next immediate time step or fixed horizon, significantly improving both the model’s utility and overall forecasting accuracy.
@@
0:  Target output length $H$, forecast horizon of each output projection $\{p_{1},p_{2},\ldots,p_{P}\}$ in descending order

### v2

@@ 1 Introduction
    We present Time-MoE, a universal decoder-only time series forecasting foundation model architecture with mixture-of-experts. To the best of our knowledge, this is the first work to scale time series foundation models up to 2.4 billion parameters. Time-MoE achieves substantial improvements in forecasting accuracy and consistently outperforms dense models with comparable computational resources, while maintaining high efficiency.
@@
    Trained on Time-300B, Time-MoE models outperform other time series foundation models with a similar number of activated parameters across six real-world benchmarks, achieving reductions in forecasting errors by an average of 20% and 24% in zero-shot and in-distribution scenarios, respectively.
@@ 3 Methodology
Problem Statement. We address the problem of predicting future values in a time series: given a sequence of historical observations $\mathbf{X}_{1:T}=\left(x_{1},x_{2},\ldots,x_{T}\right)\in\mathbb{R}^{T}$ spanning $T$ time steps, our objective is to forecast the next $H$ time steps, i.e., $\hat{\mathbf{X}}_{T+1:T+H}=f_{\theta}\left(\mathbf{X}_{1:T}\right)\in\mathbb{R}^{H}$. Here, $f_{\theta}$ represents a time series model, where $T$ is the context length and $H$ is the forecasting horizon. Notably, both $T$ and $H$ can be flexible during Time-MoE inference, distinguishing it from task-specific models with fixed horizons. Additionally, channel independence (Nie et al., 2023) is adopted to transform a multivariate input into univariate series, allowing Time-MoE to handle any-variate forecasting problems in real-world applications.
Figure 2: The architecture of Time-MoE, which is a decoder-only model. Given an input time series of arbitrary length, 1 we first tokenize it into a sequence of data points, 2 which are then encoded. These tokens are processed through $N$-stacked backbone layers, primarily consisting of causal multi-head self-attention and 3 sparse temporal mixture-of-expert layers. During training, 4 we optimize forecasting heads at multiple resolutions. For model inference, Time-MoE provides forecasts of flexible length by 5 dynamically scheduling these heads.
@@ 3.1 Time-MoE Overview
$g_{N+1,t}$ $=\operatorname{Sigmoid}\left(\mathbf{W}^{l}_{N+1}\mathbf{\bar{u}}^{l}_{t}\right),$ (7)
@@
Multi-resolution Forecasting. We introduce a novel multi-resolution forecasting head, which allows for forecasting at multiple scales simultaneously, in contrast to existing foundation models that are limited to a single fixed scale. This capability enhances Time-MoE ’s flexibility by enabling forecasting across various horizons. The model employs multiple output projections from single-layer FFNs, each designed for different prediction horizons. During training, Time-MoE aggregates forecasting errors from different horizons to compute a composite loss (Section 3.2.2), thereby improving the model generalization. By incorporating a simple greedy scheduling algorithm (see Appendix B), Time-MoE efficiently handles predictions across arbitrary horizons. This design also boosts prediction robustness through multi-resolution ensemble learning during inference.
@@ 3.2.1 Time-300B Dataset
Training time series foundation models require extensive, high-quality data. However, well-processed large-scale datasets are still relatively scarce. Recent advancements have facilitated the collection of numerous time series datasets from various sources (Godahewa et al., 2021; Ansari et al., 2024; Woo et al., 2024; Liu et al., 2024b). Nonetheless, data quality still remains a challenge, with prevalent issues such as missing values and invalid observations (Wang et al., 2024a) that can impair model performance and destabilize training. To mitigate these issues, we developed a streamlined data-cleaning pipeline (Appendix C) to filter and refine raw data, and constructed the largest open-access, high-quality time series data collection named Time-300B for foundation model pre-training. Time-300B is composed of a diverse range of publicly available datasets, spanning multiple domains such as energy, retail, healthcare, weather, finance, transportation, and web, as well as a portion of synthetic data to enhance the data quantity and diversity. Time-300B covers a wide spectrum of sampling frequencies from seconds to yearly intervals, and has over 300 billion time points after being processed by our data-cleaning pipeline, as summarized in Table 1.
@@
3.2.2 Loss Function
Pre-training time series foundation models in large scale presents significant challenges in training stability due to the massive datasets and the vast number of parameters involved. To address this, we use the Huber loss (Huber, 1992; Wen et al., 2019), which provides greater robustness to outliers and improves training stability:
where $\delta$ is a hyperparameter that balances the L1 and L2 loss components.
When training the model with a MoE architecture, focusing solely on optimizing prediction error often leads to load imbalance issues among the experts. A common problem is routing collapse (Shazeer et al., 2017), where the model predominantly selects only a few experts, limiting training opportunities for others. To mitigate this, following the approaches of (Dai et al., 2024; Fedus et al., 2022), we achieve expert-level balancing with an auxiliary loss to reduce routing collapse:
$\mathcal{L}=\frac{1}{P}\sum_{j=1}^{P}\mathcal{L}_{\text{ar}}\left(\mathbf{X}_{t+1:t+p_{j}},\hat{\mathbf{X}}_{t+1:t+p_{j}}\right)+\alpha\mathcal{L}_{\text{aux}},$ (11)
where $P$ is the number of multi-resolution projections and $p_{j}$ is the horizon of the $j$-th projection.
3.2.3 Model Configurations and Training Details
[…]
@@ 4 Main Results
Time-MoE consistently outperforms state-of-the-art forecasting models by large margins across six well-established benchmarks and settings (Appendix B). To ensure a fair comparison, we adhered to the experimental configurations from (Woo et al., 2024) for out-of-distribution forecasting and (Wu et al., 2023a) for in-distribution forecasting with a unified evaluation pipeline we developed. Specifically, we evaluate Time-MoE against 16 different baselines, representing state-of-the-art models in long-term forecasting. They are categorized into two groups: 1. zero-shot forecasting evaluation group, includes pre-trained foundation models such as Moirai (2024), TimesFM (2024), Moment (2024), and Chronos (2024); 2. in-distribution (full-shot) forecasting evaluation group, consists of modern time series models such as iTransformer (2024a), TimeMixer (2024b), TimesNet (2023a), PatchTST (2023), Crossformer (2023), TiDE (2023), DLinear (2023), and FEDformer (2022b).
@@ 4.1 Zero-shot Forecasting
Table 3: Full results of zero-shot forecasting experiments. A lower MSE or MAE indicates a better prediction. TimesFM, due to its use of Weather datasets in pretraining, is not evaluated on these two datasets and is denoted by a dash (--). Red: the best, Blue: the 2nd best.
@@
Results. Detailed results of zero-shot forecasting are in Table 3. Time-MoE achieves consistent state-of-the-art performances, improving a large margin as MSE reduction in average exceeding 20% over the other most competitive baselines. Importantly, as the model size scales (e.g., base $\rightarrow$ ultra), it continuously exhibits enhanced performance across all datasets, affirming the efficacy of scaling laws within our time series foundation models. Furthermore, in comparisons with robust baselines that have a similar number of activated parameters, Time-MoE demonstrates significantly superior performance. The largest models among the state-of-the-art baselines are Chronoslarge, Moment and Moirailarge. Compared to those models, Time-MoE achieved average MSE reductions of 23%, 30% and 11% respectively.
@@ 4.2 In-distribution Forecasting
Table 4: Full results of in-domain forecasting experiments. A lower MSE or MAE indicates a better prediction. Full-shot results besides Global Temp are obtained from (Liu et al., 2024a). Red: the best, Blue: the 2nd best.
@@
0.306 0.362 0.304 0.359 0.301 0.358 0.349 0.382 0.337 0.375 0.356 0.400 0.349 0.382 0.560 0.516 0.421 0.439 0.387 0.416 0.375 0.417
$1^{\text{st}}$ Count 4 21 33 0 7 0 0 0 0 0 0
Setup. We fine-tune the pre-trained Time-MoE models on the train split of the above-mentioned six benchmarks and set the number of finetuning epochs to only one.
Results. The full results are in Table 4. Time-MoE exhibits remarkable capabilities, comprehensively surpassing advanced deep time series models from recent years, achieving a MSE reduction of 24% in average. Fine-tuning on downstream data with only one epoch significantly improves predictive performance, showcasing the remarkable potential of large time series models built on the MoE architecture. Similar to zero-shot forecasting, as the model size increases, the scaling law continues to be effective, leading to continuous improvements in the performance of the Time-MoE.
@@ 4.3 Ablation Study
Time-MoEbase 0.262
         w/o Huber loss 0.267
         w/o multi-resolution layer 0.269
         w/o mixture-of-experts 0.272
         w/o auxiliary loss 0.275
@@
Time-MoEbase 0.262 0.095 s/iter
Time-MoEbase w/ {1,8,32} 0.273 0.130 s/iter
Time-MoEbase w/ {1,8} 0.320 0.411 s/iter
Time-MoEbase w/ {1} 1.382 2.834 s/iter
@@
Replacing the MoE layers with standard FFNs (w/o mixture-of-experts) led to an average performance drop from $0.262$ to $0.272$, highlighting the performance boost provided by the sparse architecture. A detailed comparison of dense and sparse models is presented in Section  4.4. We retained only the horizon-32 output layer by eliminating the other multi-resolution output layers from the Time-MoEbase, excluding the multi-task optimization (w/o multi-resolution layer). Consequently, we observed that the performance of this modified model was slightly inferior compared to that of the Time-MoEbase. Additionally, as shown in the right side of Table 5, our default selection of four multi-resolution output projections with receptive horizons of $\{1,8,32,64\}$ results in optimal predictive performance and inference speed. As we reduce the number of multi-resolution output projections, performance consistently declines, and inference speed significantly increases. This demonstrates the rationality of our multi-resolution output projection design.
@@ 4.4 Scalability Analysis
. Average MSE Training Speed Inference Speed Training Memory Inference Memory Time-MoEbase 0.262 0.84 s/iter 0.095 s/iter 1.77 GB 226.70 MB Time-MoEbase w/o FA 0.262 1.09 s/iter 0.118 s/iter 1.77 GB 226.70 MB Time-MoEbase w/ FP32 0.261 1.24 s/iter 0.133 s/iter 2.21 GB 453.41 MB
@@ 4.5 Sparsification Analysis
w/ {Top1} 0.264 0.082 s/iter
w/ {Top2} 0.262 0.095 s/iter
w/ {Top4} 0.262 0.109 s/iter
w/ {Top6} 0.265 0.120 s/iter
w/ {Top8} 0.269 0.129 s/iter
We performed a sensitivity analysis on the number of experts, represented as topk, within the Time-MoE architecture, as shown in Table 7. As $k$ increases, performance shows only marginal changes, with minimal improvements in average MSE. However, inference time increases noticeably as more experts are utilized. This indicates that increasing sparsity within the MoE architecture does not compromise performance but significantly enhances computational efficiency. This balance is critical for scaling time series foundation models, where optimizing performance and computational cost is essential. Sparse MoE architectures inherently offer advantages in these areas.
@@ 5 Conclusion
In this paper, we introduced Time-MoE, a scalable and unified architecture for time series foundation models that leverages a sparse design with mixture-of-experts to enhance computational efficiency without compromising model capacity. Pre-trained on our newly introduced large-scale time series dataset, Time-300B, Time-MoE was scaled to 2.4 billion parameters, with 1.1 billion activated, demonstrating significant improvements in forecasting accuracy. Our results validate the scaling properties in time series forecasting, showing that Time-MoE consistently outperforms dense models with equivalent computational budgets across multiple benchmarks. With its ability to perform universal forecasting and superior performance in both zero-shot and fine-tuned scenarios, Time-MoE establishes itself as a state-of-the-art solution for real-world forecasting challenges. This work paves the way for future advancements in scaling and enhancing the efficiency of time series foundation models.
@@ Appendix B Implementation Details
We evaluate the performance of various models for long-term forecasting across eight well-established datasets, including the Weather (Wu et al., 2021), Global Temp (Wu et al., 2023b), and ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2) (Zhou et al., 2021). A detailed description of each dataset is provided in Table 9.
@@
To construct the multi-resolution forecasting head, we define $P$ output projections, each corresponding to a distinct forecasting horizon, denoted as $\left(p_{1},p_{2},\ldots,p_{P}\right)$. The output projection for horizon $p_{j}$ is used to forecast the subsequent $p_{j}$ time steps, as follows:
$\hat{\mathbf{X}}_{t+1:t+p_{j}}=\mathbf{W}_{p_{j}}\mathbf{h}^{L}_{t},$ (12)
where $\mathbf{W}_{p_{j}}\in\mathbb{R}^{p_{j}\times D}$ is the learnable parameter matrix for that horizon, and $\mathbf{h}^{L}_{t}$ represents the output hidden state from the last MoE Transformer block. All output projections are optimized simultaneously during model training.
During inference, we apply a greedy scheduling algorithm for arbitrary target output lengths $H$, as outlined in Algorithm 1. For each forecast operation in the auto-regressive process, we select a projection $p_{j}$ with the closest forecasting horizon that does not exceed the remaining forecast duration. This approach allows Time-MoE to extend predictions beyond the next immediate time step or fixed horizon, significantly improving both the model’s utility and overall forecasting accuracy.
@@
0:  Target output length $H$, forecast horizon of each output projection $\{p_{1},p_{2},\ldots,p_{P}\}$ in ascending order